Automated Derivation of Limiting Distributions Of Combinatorial Random Variables given by Bivariate Rational Functions By Gleb Pogudin, Manuel Kauers, and Doron ZEILBERGER

Abstract:

0. Maple Package and Sample Output This article is accompanied by a Maple package, BVM.txt, that is available, along with six input and output files, from

http://www.math.rutgers.edu/~zeilberg/mamarim/mamarimhtml/bvm.html .

1. Preface

8. References

Gleb Pogudin, Institute for Algebra, Johannes Kepler Universty, Linz, Austria. pogudin.gleb@gmail.com

Manuel Kauers, Manuel Kauers ,Institute for Algebra, Johannes Kepler Universty, Linz, Austria.

manuel.kauers@jku.at

Doron Zeilberger, Department of Mathematics, Rutgers University (New Brunswick), Hill Center-Busch Campus, 110 Frelinghuysen Rd., Piscataway, NJ 08854-8019, USA;

DoronZeil at gmail dot com ; http://www.math.rutgers.edu/~zeilberg/ .

Exclusively published in The Personal Journal of Shalosh B. Ekhad and Doron Zeilberger

(http://www.math.rutgers.edu/~zeilberg/pj.html), the authors’ websites, and arxiv.org .

Dec. 15, 2016